One wall of a room is 8 feet by 10 feet, and a rectangular door 3 feet by 7 feet is to be cut in the wall. The remainder of the wall is to be wallpapered. There are three other walls in the room that are 8 feet by 10 feet and these are also to be wallpapered. What is the total area that is to be wallpapered?
Thank you

Answer by mananth(16949)

Total area of all the walls less area of door = wallpaper area
Area of one wall = 8*10 = 80 sq.ft
area of 4 walls = 89*4 = 320 sq.ft
area of door = 7*3 =21 sq.ft
wallpaper area = 320-21= 299 sq. ft